FUL LSCREEN PRINTER VERSION ACK Chapter 21, Problem 027 The magnltude of the electrostatic frce betvren two identical ions that are separated by a distance of 8.1 x 1010 m is 406.6x 10 N. Co) What is the charge of eadh ion? (b) How many electrons are missing from each ion (thus giving the ion its charge imbalance)? (a) Number (b) Number Units Units Question Attempts: unlimited Dro コ SAVE FOR LATER SUBMET ANSWER
Answer #1

Electrostatic Force ,F=Kq_{1}q_{2}/d^{2}

here charges are identical

soF=Kq^{2}/d^{2}

406.6 10N9 10/(8.1 10-10m)2

406.6 109/(8.1* 10-10)-2

406.6 * 10-9 1.372 * 1028 * q2

= 2.964 * 10-35

ANSWER: = 5.44 * 10-18C

==========================

q =charge of eletron *no .of electron

q=e *n

18 5.44 10 5C-1.60210n

18 5.44 * 10- /(1.602 * 10-19) = n

33.9575

ANSWER: 34

================